A solution of 30% acid and a solution of 60% acid are to be mixed to produce 50 liters of 57% acid. How many liters of each solution should be used?

Answer by checkley79(3341)

.60x+.30(50-x)=.57*50
.60x+15-.30x=28.5
.30x=28.5-15
.30x=13.5
x=13.5/.30
x=45 liters of 60% solution is used.
50-45=5 liters of 30% solution is used.
Proof:
.60*45+.30*5=.57*50
27+1.5=28.5
28.5=28.5